# You Only Pose Once: A Minimalist’s Detection Transformer for Monocular RGB Category-level 9D Multi-Object Pose Estimation

### v2

@@ Abstract
Accurately recovering the full 9-DoF pose of unseen instances within specific categories from a single RGB image remains a core challenge for robotics and automation. Most existing solutions still rely on pseudo-depth, CAD models, or multi-stage cascades that separate 2D detection from pose estimation. Motivated by the need for a simpler, RGB-only alternative that learns directly at the category level, we revisit a longstanding question: Can object detection and 9-DoF pose estimation be unified with high performance, without any additional data? We show that they can be achieved with our method, YOPO, a single-stage, query-based framework that treats category-level 9-DoF estimation as a natural extension of 2D detection. YOPO augments a transformer detector with a lightweight pose head, a bounding-box–conditioned translation module, and a 6D-aware Hungarian matching cost. The model is trained end-to-end only with RGB images and category-level pose labels. Despite its minimalist design, YOPO sets a new state of the art on three benchmarks. On the REAL275 dataset, it achieves 79.6% IoU50 and 54.1% under the $10^{\circ}$$10{\rm{cm}}$ metric, surpassing all prior RGB-only methods and closing much of the gap to RGB-D systems. The code, models, and additional qualitative results can be found on our project page [^1] https://mikigom.github.io/YOPO-project-page.
@@ III-A Overall Architecture
In the first stage, the transformer encoder processes the multi-scale backbone features, and a detection head predicts a set of reference points with associated objectness scores. The top-$K$ proposals are selected based on these scores and are transformed into object queries $\mathbf{q}$, which carry both spatial (2D reference point) and semantic information. In the second stage, the transformer decoder refines these queries to produce $\mathbf{p}$, which are subsequently fed into parallel prediction heads for classification, 2D box regression, and 9D pose estimation.
@@ III-B Parallel Prediction Heads
Pose Estimation Head. The pose-estimation head predicts the 9D pose parameters $(\mathbf{R}_{i},\mathbf{t}_{i},\mathbf{s}_{i})$ via four regression branches: (1) 2D center offset, (2) depth, (3) rotation, and (4) scale. This head is supervised in both stages: during the proposal stage, auxiliary supervision encourages object queries to encode early geometric information; during the refinement stage, it outputs the final 9D pose estimates. The 3D translation $\mathbf{t}_{i}$ is reconstructed by back-projecting the predicted 2D center and depth using the camera intrinsics $K$.
This parallel design allows the detection head to specialize in object localization, while the pose-estimation head focuses on 3D reasoning. Early supervision of the pose head ensures that object queries are geometry-aware even before refinement. By sharing the same set of object queries, our architecture jointly optimizes detection and pose estimation, enabling the two tasks to reinforce each other during training.
@@
For 3D translation, we employ a disentangled parameterization following Simonelli et al. [41], wherein the image-plane center and depth are predicted separately to enhance training stability in monocular settings. Specifically, the 2D center $\mathbf{u}_{i}$ is normalized by the image width and height so that it lies in $0,1]^{2}$, and a sigmoid activation is applied to the corresponding outputs of the linear head. During inference, these normalized coordinates are rescaled by the original image dimensions to recover pixel-space positions. The final 3D translation $\mathbf{t}_{i}$ is reconstructed by back-projecting the rescaled 2D center and the predicted depth using the camera intrinsic matrix $K$.
Bounding Box-Conditioned Center Prediction. Simonelli et al. [41] proposed regressing the 2D center $\mathbf{u}_{i}$ as an offset from the center of the predicted 2D bounding box:
$\mathbf{u}_{i}$ $=\mathbf{b}_{i}^{\text{center}}+\Delta\mathbf{u}_{i},\quad\Delta\mathbf{u}_{i}=\text{MLP}(\mathbf{p}_{i}),$ (2)
where $\mathbf{p}_{i}\in\mathbb{R}^{D}$ is the refined embedding of the $i$-th object, obtained from the transformer decoder output, and $\mathbf{b}_{i}^{\text{center}}$ denotes the geometric center of the predicted bounding box $\mathbf{b}_{i}$. MLP refers to the linear head.
We further enhance the center-offset prediction head by explicitly conditioning it on the bounding-box parameters, as shown in Figure 2(b). Specifically, instead of relying solely on the object query $\mathbf{p}_{i}$, we concatenate it with the predicted bounding box $\mathbf{b}_{i}\in\mathbb{R}^{4}$ to form an augmented input to the MLP:
$\Delta\mathbf{u}_{i}$ $=\operatorname{MLP}\!\big(\operatorname{Concat}(\mathbf{p}_{i},\mathbf{b}_{i})\big),\quad\operatorname{MLP}:\ \mathbb{R}^{D+4}\rightarrow\mathbb{R}^{2},$
$\mathbf{u}_{i}$ $=\mathbf{b}_{i}^{\text{center}}+\Delta\mathbf{u}_{i}.$ (3)
This design allows the offset prediction to be informed not only by semantic features but also by explicit geometric guidance encoded in the bounding box.
At inference time, the predicted 2D center and depth are combined through perspective back-projection to yield the final 3D object translation $\mathbf{t}_{i}$ in camera coordinates:
$\mathbf{t}_{i}=z_{i}\,K^{-1}\begin{bmatrix}\mathbf{u}_{i}\\ 1\end{bmatrix}$. (4)
This formulation maintains the structural benefits of the disentangled translation framework while enhancing its representational power through bounding box-conditioned center prediction.
Bounding Box-Conditioned Depth Prediction. Following the center head, we extend the depth predictor by concatenating the object query with the bounding-box parameters. This conditioning provides complementary geometric cues, enabling more stable depth regression and reducing scale ambiguities, ultimately improving translation accuracy.
@@
where $\mathcal{C}_{\text{cls}}$ is the classification cost, $\mathcal{C}_{\text{bbox}}$ is the $\ell_{1}$ distance between the predicted and ground-truth 2D bounding boxes, $\mathcal{C}_{\text{IoU}}$ is the negative IoU, and $\lambda^{\mathcal{C}}$ denotes the corresponding weights.
To handle 6D pose estimation, we add translation and rotation terms: $\mathcal{C}_{\text{trans}}$ is the Euclidean distance between 3D translations, and $\mathcal{C}_{\text{rot}}$ is the geodesic distance between rotation matrices, optionally accounting for object symmetries. The final matching cost is
@@
We use the following default weights: $\lambda^{\mathcal{C}}_{\text{cls}}{=}2.0$, $\lambda^{\mathcal{C}}_{\text{bbox}}{=}5.0$, $\lambda^{\mathcal{C}}_{\text{IoU}}{=}2.0$, $\lambda^{\mathcal{C}}_{\text{trans}}{=}5.0$, $\lambda^{\mathcal{C}}_{\text{rot}}{=}2.0$.
@@ III-E Implementation Details
Pose Estimation Head. We represent rotations using the continuous 6D parameterization [42] and supervise them with a geodesic loss on $SO(3)$. Both the depth $z$ and the anisotropic 3D scale $\mathbf{s}\in\mathbb{R}^{3}$ are regressed directly in linear space. Unless otherwise stated, we adopt class-wise rotation and scale prediction heads: the output dimensions are expanded by the number of categories, and the final predictions are selected according to the highest class confidence.
@@ IV-C Ablation Study and Discussion
We ablate YOPO on REAL275 with emphasis on ${\rm IoU}_{50}$ and the $10^{\circ}10\text{cm}$ metric. Starting from a model with all components disabled, performance is modest (51.2 ${\rm IoU}_{50}$ / 26.1 $10^{\circ}10\text{cm}$). Introducing box-conditioning on the center head (BC {center}) yields a sizable overlap gain (59.8 ${\rm IoU}_{50}$; +8.6), albeit at the cost of geometric accuracy (21.8 $10^{\circ}10\text{cm}$). Adding data augmentation substantially recovers the latter (34.1 $10^{\circ}10\text{cm}$) while keeping overlap stable (59.1 ${\rm IoU}_{50}$). With augmentation but without class-wise heads, proper weight scaling (WS) within 3D-aware matching raises performance to 66.6 ${\rm IoU}_{50}$ and 42.4 $10^{\circ}10\text{cm}$. Extending BC to also condition the depth head (BC {center, z}) provides a consistent overlap gain (67.1 ${\rm IoU}_{50}$) with comparable $10^{\circ}10\text{cm}$ (40.7). Further gains come from fine-tuning on REAL275 (RFT), which raises the ResNet-50 model to 71.1 on ${\rm IoU}_{50}$ and 46.6 on $10^{\circ}10\text{cm}$; with a stronger Swin-L backbone, the same recipe reaches 79.6 on ${\rm IoU}_{50}$ and 54.1 on $10^{\circ}10\text{cm}$. Overall, most of the performance improvement is driven by 3D-aware matching, box-conditioning, and supporting components. The shaded row indicates our final model, which offers a balanced and strong performance.
@@
Table V isolates the effect of conditioning different prediction heads on the 2D box. Without any conditioning, the model attains 66.0 on ${\rm IoU}_{50}$ and 39.5 on $10^{\circ}10\text{cm}$. Conditioning only the center slightly degrades the main metrics (64.2 on ${\rm IoU}_{50}$ / 37.4 on $10^{\circ}10\text{cm}$), and adding conditioning to rotation or size (on top of the center) yields limited or inconsistent gains. In contrast, conditioning depth together with the center produces a clear joint improvement: 67.1 on ${\rm IoU}_{50}$ and 40.7 on $10^{\circ}10\text{cm}$. While extending conditioning further can improve a single metric, it comes with trade-offs on the complementary objective. Therefore, we adopt conditioning only the center and depth heads as the default, as this choice offers the best balance between overlap and geometric accuracy.
@@
In Table VI, we examine whether our method requires precise 2D bounding boxes during training. Specifically, we compare the standard setup that uses directly annotated 2D boxes () with a weaker alternative in which 2D boxes are derived from projected 3D cuboids (), which are less accurate. On ResNet-50, precise 2D boxes yield slightly better 3D IoU (+1.2 points on ${\rm IoU}_{75}$) and translation accuracy (+2.8 points on the $10\,\text{cm}$ metric), whereas the projected-cuboid boxes still produce competitive results and even achieve higher rotation accuracy (+3.3 points on the $10^{\circ}$ metric). With the Swin-L backbone, both strategies deliver strong performance, indicating that our method remains robust even under weaker box supervision. Overall, YOPO does not strictly rely on highly accurate 2D bounding boxes, which can further reduce annotation costs in practical settings.
@@
Table VII compares YOPO (Swin-L) against RGB-D and RGB methods on REAL275 and HouseCat6D. On REAL275, YOPO decisively surpasses RGB-only baselines in both overlap and geometric accuracy, approaching RGB-D performance in overlap while trailing on stricter criteria (e.g., ${\rm IoU}_{75}$ and $10^{\circ}5\text{cm}$). On HouseCat6D, YOPO achieves 34.8 on ${\rm IoU}_{50}$ and 5.3 on $5^{\circ}5\text{cm}$, delivering performance comparable to GPV-Pose on the latter (4.6) and exceeding NOCS in ${\rm IoU}_{50}$ (21.2), though it remains behind the strongest RGB-D system (SpotPose). Notably, methods reported on HouseCat6D leverage ground-truth segmentation masks, whereas YOPO operates purely on RGB images without such auxiliary supervision. Overall, YOPO markedly narrows the gap between RGB and RGB-D methods and establishes a clear lead among RGB-only approaches, underscoring the effectiveness of our end-to-end framework for category-level 9D pose estimation from RGB.
Inference Time
Our model performs joint object detection and 9D pose estimation for multiple objects in a single RGB image in a single forward pass. It achieves 21.3 frames per second (FPS) with ResNet-50 and 7.7 FPS with Swin-Large on an NVIDIA A6000 GPU, combining state-of-the-art accuracy with efficient end-to-end inference.
@@ Acknowledgement
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2025-02653113, High-Performance Research AI Computing Infrastructure Support at the 2 PFLOPS Scale). This research was supported by Deep-Tech Tips funded by Ministry of SMEs and Startups.

### v3

@@ Abstract
Accurately recovering the full 9-DoF pose of unseen instances within specific categories from a single RGB image remains a core challenge for robotics and automation. Most existing solutions still rely on pseudo-depth, CAD models, or multi-stage cascades that separate 2D detection from pose estimation. Motivated by the need for a simpler, RGB-only alternative that learns directly at the category level, we revisit a longstanding question: Can object detection and 9-DoF pose estimation be unified with high performance, without any additional data? We show that they can be achieved with our method, YOPO, a single-stage, query-based framework that treats category-level 9-DoF estimation as a natural extension of 2D detection. YOPO augments a transformer detector with a lightweight pose head, a bounding-box–conditioned translation module, and a 6D-aware Hungarian matching cost. The model is trained end-to-end only with RGB images and category-level pose labels. Despite its minimalist design, YOPO sets a new state of the art on three benchmarks. On the REAL275 dataset, it achieves 79.6% IoU50 and 54.1% under the $10^{\circ}$$10{\rm{cm}}$ metric, surpassing all prior RGB-only methods and closing much of the gap to RGB-D systems. The code, models, and additional qualitative results can be found on our project page [^1] https://mikigom.github.io/YOPO-project-page/.
@@ III-A Overall Architecture
In the first stage, the transformer encoder takes the flattened multi-scale backbone features as input and outputs memory features enriched with global context. A detection head then predicts a set of reference points with associated objectness scores. The top-K proposals are selected based on these scores and are used as spatial anchors (2D reference points) with learnable object queries $\mathbf{q}$. In the second stage, the transformer decoder takes these queries and iteratively updates them through cross-attention with the encoder’s memory features to produce enriched embeddings $\mathbf{p}$. These refined queries are subsequently fed into parallel prediction heads to explicitly derive the 9-DoF parameters: rotation ($\mathbf{R}_{i}$), translation ($\mathbf{t}_{i}$), and scale ($\mathbf{s}_{i}$).
@@ III-B Parallel Prediction Heads
Pose Estimation Head. The pose estimation head takes the object queries and predicts the 9-DoF parameters $(\mathbf{R}_{i},\mathbf{t}_{i},\mathbf{s}_{i})$ via four specialized multi-layer perceptron (MLP) branches: (1) 2D center offset, (2) depth, (3) rotation, and (4) scale. Specifically, the scale head directly regresses the 3D anisotropic scale $\mathbf{s}_{i}$. By leveraging the global context encoded in the query, it predicts absolute scale without relying on categorical shape priors. To alleviate monocular ambiguity, both the center and depth heads predict their values by explicitly conditioning the query on the 2D bounding box (detailed in the following subsection). Rotation is predicted via a continuous 6D representation. The 3D translation $\mathbf{t}_{i}$ is reconstructed by back-projecting the predicted 2D center and depth using the camera intrinsics $K$. This head is supervised at both the proposal and refinement stages, ensuring queries are geometry-aware early on.
By sharing object queries, our parallel architecture jointly optimizes 2D localization and 3D reasoning, enabling the two tasks to mutually reinforce each other.
@@
For 3D translation, we employ a disentangled parameterization following Simonelli et al. [40], wherein the image-plane projected center $(u_{i},v_{i})$ and physical depth $z_{i}$ are predicted separately. This decoupled approach significantly enhances training stability in monocular settings.
Bounding Box-Conditioned Center Prediction. Following the disentangled translation formulation of Simonelli et al. [40], we predict the image-plane center as an offset from the center of the predicted 2D bounding box. Let $\mathbf{p}_{i}\in\mathbb{R}^{D}$ denote the refined embedding of the $i$-th object from the transformer decoder, and $\mathbf{b}_{i}\in\mathbb{R}^{4}$ the predicted bounding box parameterized as $\mathbf{b}_{i}=(c_{x,i},c_{y,i},w_{i},h_{i})$. The first two components $\mathbf{b}_{i,1:2}\in\mathbb{R}^{2}$ correspond to the box center.
Instead of regressing the center solely from the object query, we condition the prediction on the bounding box by concatenating $\mathbf{p}_{i}$ and $\mathbf{b}_{i}$:
$\mathbf{c}_{raw,i}=\operatorname{MLP}_{c}\big(\operatorname{Concat}(\mathbf{p}_{i},\mathbf{b}_{i})\big),\quad\operatorname{MLP}_{c}:\mathbb{R}^{D+4}\rightarrow\mathbb{R}^{2}.$ (2)
The predicted residual is combined with the box center and converted to a normalized center coordinate via
$\mathbf{c}_{norm,i}=\sigma(\mathbf{c}_{raw,i})+\mathbf{b}_{i,1:2}-\begin{bmatrix}0.5\\ 0.5\end{bmatrix},\quad\mathbf{c}_{norm,i}\in\mathbb{R}^{2},$ (3)
where $\sigma(\cdot)$ denotes the element-wise sigmoid. Crucially, the model is supervised directly on these normalized coordinates $\mathbf{c}_{norm,i}\in[0,1]^{2}$ during training, which ensures stability across varying image resolutions.
During inference, the normalized center is scaled by the image width $W$ and height $H$ to recover the absolute pixel coordinates:
$\begin{split}u_{i}&=\mathrm{clamp}(W\cdot\mathbf{c}_{norm,i,1},0,W),\\ v_{i}&=\mathrm{clamp}(H\cdot\mathbf{c}_{norm,i,2},0,H).\end{split}$ (4)
Finally, the 3D translation is recovered by perspective back-projection using the predicted depth $z_{i}$:
$\mathbf{t}_{i}=z_{i}K^{-1}\begin{bmatrix}u_{i}\\ v_{i}\\ 1\end{bmatrix}$. (5)
Conditioning the center prediction on the bounding box provides explicit geometric guidance, allowing the model to exploit spatial cues encoded in the detection head while preserving the structural benefits of the disentangled translation formulation.
Bounding Box-Conditioned Depth Prediction. Similar to the center prediction head, we condition the depth predictor by concatenating the object query with the 2D bounding-box parameters. This explicit geometric conditioning provides complementary spatial cues, which stabilizes depth regression, reduces scale ambiguities, and ultimately improves 3D translation accuracy.
@@
where $\mathcal{C}_{\text{cls}}$ is the classification cost, $\mathcal{C}_{\text{bbox}}$ is the L1 distance between the predicted and ground-truth 2D bounding boxes, $\mathcal{C}_{\text{IoU}}$ is the negative IoU, and $\lambda^{\mathcal{C}}$ denotes the corresponding weights.
While standard DETR relies solely on 2D proximity, we explicitly regularize 3D structure by adding translation and rotation terms: $\mathcal{C}_{\text{trans}}$ is the Euclidean distance between 3D translations, and $\mathcal{C}_{\text{rot}}$ is the geodesic distance between rotation matrices (accounting for object symmetries). The final 6D-aware matching cost is
@@
We use the following default weights: $\lambda^{\mathcal{C}}_{\text{cls}}{=}2.0$, $\lambda^{\mathcal{C}}_{\text{bbox}}{=}5.0$, $\lambda^{\mathcal{C}}_{\text{IoU}}{=}2.0$, $\lambda^{\mathcal{C}}_{\text{trans}}{=}5.0$, $\lambda^{\mathcal{C}}_{\text{rot}}{=}2.0$. We omit 3D scale from $\mathcal{C}_{\text{match}}$. The base cost ($\mathcal{C}_{\text{bbox}}$ and $\mathcal{C}_{\text{IoU}}$) already provides a strong proxy for apparent object size. Furthermore, explicit 3D scale prediction from a monocular image is inherently ambiguous, especially early in training. Including it in the cost matrix can introduce noisy query assignments. Thus, we decouple scale from the matching process and optimize it purely via the loss function post-assignment.
@@ III-E Implementation Details
Pose Estimation Head. We represent rotations using the continuous 6D parameterization [41], mapped to $\mathrm{SO}(3)$ via Gram–Schmidt-like orthonormalization, and supervise them with a geodesic loss. Both the depth $z_{i}$ and the anisotropic 3D scale $\mathbf{s}_{i}\in\mathbb{R}^{3}$ are regressed directly in linear space. Unless otherwise stated, we adopt class-wise (CW) rotation and scale prediction heads: the output dimensions are expanded by the number of categories, and the final predictions are selected according to the highest class confidence.
@@ IV-C Ablation Study and Discussion
Table III details the contribution of each component. The baseline yields modest results. Bounding-box conditioning on the center head (BC {center}) improves overlap, while data augmentation (Aug) recovers geometric accuracy. Applying proper weight scaling (WS) without class-wise (CW) heads significantly boosts performance. Subsequently, enabling CW heads and extending conditioning to the depth head (BC {center, z}) achieves an optimal balance and consistent overlap gains. Finally, fine-tuning on REAL275 (RFT) pushes performance to its peak, yielding our final configuration (shaded rows).
@@
Table V shows the effect of conditioning different prediction heads on the 2D box. Without any conditioning, the model attains 66.0 on ${\rm IoU}_{50}$ and 39.5 on $10^{\circ}10\text{cm}$. Conditioning only the center slightly degrades the main metrics (64.2 on ${\rm IoU}_{50}$ / 37.4 on $10^{\circ}10\text{cm}$), and adding conditioning to rotation or size (on top of the center) yields limited or inconsistent gains. In contrast, conditioning depth together with the center produces a clear joint improvement: 67.1 on ${\rm IoU}_{50}$ and 40.7 on $10^{\circ}10\text{cm}$. While extending conditioning further can improve a single metric, it comes with trade-offs on the complementary objective. Therefore, we adopt conditioning only the center and depth heads as the default, as this choice offers the best balance between overlap and geometric accuracy.
@@
Table VI compares training with directly annotated 2D boxes () against weaker boxes derived from projected 3D cuboids (). On ResNet-50, precise boxes slightly improve overlap (+1.2 on ${\rm IoU}_{75}$) and translation (+2.8 on $10\,\text{cm}$), whereas projected boxes yield better rotation (+3.3 on $10^{\circ}$). With Swin-L, both strategies perform similarly. This indicates YOPO is robust to weaker box supervision and does not strictly rely on highly accurate 2D annotations, which can effectively reduce annotation costs in practice.
@@
Table VII compares YOPO against RGB and RGB-D methods. On REAL275, YOPO significantly outperforms all RGB-only baselines and approaches RGB-D overlap performance, though it trails on stricter metrics (e.g., ${\rm IoU}_{75}$, $10^{\circ}5\text{cm}$). On HouseCat6D, YOPO achieves 34.8 on ${\rm IoU}_{50}$ and 5.3 on $5^{\circ}5\text{cm}$, outperforming NOCS and rivaling GPV-Pose. Notably, YOPO achieves this without the ground-truth segmentation masks required by the RGB-D baselines. Overall, our method establishes a clear state-of-the-art for RGB-only approaches and substantially narrows the gap to RGB-D systems.
Inference Time & Bottleneck Analysis
Our model performs joint object detection and 9D pose estimation in a single forward pass. On an RTX A6000 GPU, it achieves $\sim$20 FPS (49.7 ms/img) with ResNet-50 and $\sim$8 FPS (124.5 ms/img) with Swin-Large. Breaking down the latency, our proposed pose head is highly lightweight, consuming only $\sim$9.1 ms regardless of the backbone. For ResNet-50, the remaining runtime consists of feature extraction (7.4 ms) and transformer processing (33.1 ms), making the transformer the primary bottleneck (67% of total latency). When scaling up to Swin-Large, heavy feature extraction (49.1 ms, 39%) and transformer processing (66.1 ms, 53%) become the main computational bottlenecks.
@@ Acknowledgement
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2025-02653113, High-Performance Research AI Computing Infrastructure Support at the 2 PFLOPS Scale) and by Deep-Tech Tips funded by the Ministry of SMEs and Startups. We also acknowledge Gyeonggi-do for providing high-performance computing resources for this research.